After an initially neutral glass rod is rubbed with an initially neutral silk scarf, the rod has a charge of 16.8 pC. Estimate the fractional increase or decrease in the scarfs mass. (Assume the scarf had a mass of 100 g) AM Would this change in mass be easily noticed? (Assume that this change is noticeable if it is more than 1 mg.) O Yes O No
Answer #1

Charge appearing on rod and scarf is due transfer of electrons.

Number of electrons transferred= charge appearing on rod/amount of charge on electron

= 16.8×10^-6(C) / 1.6×10^-19(C) = 1.05×10^14

Transfer of mass or changr in mass =

Number of electrons transferred*mass of electron

= 1.05×10^14 * 9.1×10^-31(kg)

= 9.5×10^-17 kg = 9.5×10-17(kg)*10^6(mg/kg)

= 9.5×10-11 mg < 1 mg

Delta M/M = 9.5×10^-17(kg)/0.1(kg)= 9.5×10^-16

Change in mass is not easily noticed as it is less than 1 mg